I am currently learning about conic sections. What would the graph of 3x^2=8-4y^2-8y and 5y^2=10-4x^2 be? circle, ellipse, hyperbola, parabola Answer by stanbon(75887) 
 What would the graph of
---
3x^2=8-4y^2-8y
---
Rearrange:
3x^2+4y^2+8y = 8
----
Since the coefficients of x^2 and y^2 have
the same sign but are unequal, the graph
would be an ellipse.
-----------------------------
5y^2=10-4x^2 be?
Rearrange:
4x^2 + 5y^2 = 10
Same reason as above.
==========================
So, ellipse.
Cheers,
Stan H.